A bank account starts with $700 and earns an annual interest rate of 6%. Use the continuous compounding formula to estimate the balance after 4 years.
B=

Answer by Theo(13342)

the continuous compounding formula is f = p * e ^ (r * t).

f is the future value.
p is the present value.
r is the interest rate per time period.
t is the number of time periods.
e is the scientific constant of 2.718281828.

in your problem, .....

p is 700.
r = 6% = .06 per year.
t = 4 yers.

the formula becomes f = 700 * e ^ (.06 * 4).

solve for f to get f = 889.8744052.

if you don't have the constant e in your calculator, then f = 700 * 2.718281828 ^ (.06 * 4) = the same.